OpenStudy (anonymous):

I am not sure how to do this equation. HELP!!! |x-1|=5x=10

hartnn (hartnn):

correct the question pleae, there are 2 = signs .... :))

OpenStudy (anonymous):

|x-1|=5x+10

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

when |a|= b then a=b or a=-b so what about |x-1|=5x+10

OpenStudy (anonymous):

How do I set it up? x-1=5x+10 or x-1=5x-10

OpenStudy (anonymous):

How do I set it up? x-1=5x+10 or x-1=5x-10

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

oh, sorry for late replies ... |x-1|=5x+10 ---> x-1 = (5x+10) or x-1 = -(5x+10) got this ?

OpenStudy (anonymous):

@hartnn yes! Thank you for helping!

hartnn (hartnn):

now can you solve those 2 equations separately ? (thank me when you get the correct final answer by yourself :) )

hartnn (hartnn):

x-1 = (5x+10) x=... ?

OpenStudy (anonymous):

x-1=(5x+10) or x-1=-(5x+10) x-1=5x+10 x-1=-5x-10 +1 +1 +1 +1 x=5x+11 x=-5x-9 -5x -5x +5x +5x -5x=11 6x=-9 x=-2.2 x=-1.5

OpenStudy (anonymous):

x-1=(5x+10) or x-1=-(5x+10) x-1=5x+10 x-1=-5x-10 +1 +1 +1 +1 x=5x+11 x=-5x-9 -5x -5x +5x +5x -5x=11 6x=-9 x=-2.2 x=-1.5

hartnn (hartnn):

oh wow! you are already good at this! :) now last step is to verify whether these 2 are actually the solutions to original equations.

hartnn (hartnn):

so put x= -2.2 in |x-1|=5x=10 does left side = right side ?

OpenStudy (anonymous):

(-1.5)-1=5(-1.5)+10 (-2.2)-1+5(-2.2)+10 |-2.5|=2.5 yes |-3.2|=-1 no

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

oh, yes , you are very correct :) so, the only solution is x = -1.5

OpenStudy (anonymous):

@hartnn thank you very much! You have been a huge help with this question!

OpenStudy (anonymous):

@hartnn thank you very much! You have been a huge help with this question!

hartnn (hartnn):

hey, you are most welcome ^_^